OpenStudy (anonymous):

Find all solutions if 0

OpenStudy (dumbcow):

oh oops, i thought those 2's were exponents

myininaya (myininaya):

are they suppose to be dola?

OpenStudy (anonymous):

Nope they are not exponents

myininaya (myininaya):

we can factor sinx out

OpenStudy (dumbcow):

no wonder i was having trouble then...:)

OpenStudy (anonymous):

:0P But still, thank you for trying dumbcow

myininaya (myininaya):

oops i made a mistake above

OpenStudy (dumbcow):

myininaya, use sum of angle formula here sin(3x) = sin(2x)cos(x) + sin(x)cos(2x)

OpenStudy (dumbcow):

sin(3x) = 1/2 sin is 1/2 at pi/6 and 5pi/6 3x = pi/6 --> x = pi/18 3x = 5pi/6 --> x = 5pi/18

myininaya (myininaya):

cool!

myininaya (myininaya):

i was thinking way too hard

myininaya (myininaya):

gj

OpenStudy (dumbcow):

yeah trig can be tricky, i didn't see it right away either

myininaya (myininaya):

its past my bedtime

OpenStudy (anonymous):

Wow! That made things so much different! haha

OpenStudy (dumbcow):

x2dola, that make sense that should complete the problem

OpenStudy (anonymous):

Yes, thank you very much!